# Uncertainty-Aware Knowledge Tracing Models

### v1

@@ Abstract
The main focus of research on Knowledge Tracing (KT) models is on model developments with the aim of improving predictive accuracy. Most of these models make the most incorrect predictions when students choose a distractor, leading to student errors going undetected. We present an approach to add new capabilities to KT models by capturing predictive uncertainty and demonstrate that a larger predictive uncertainty aligns with model incorrect predictions. We show that uncertainty in KT models is informative and that this signal would be pedagogically useful for application in an educational learning platform that can be used in a limited resource setting where understanding student ability is necessary.
@@ 1 Introduction
Knowledge Tracing (KT) models are widely used to predict a student’s response to a future question given their interaction history. KT models have developed over time from earlier work on Bayesian KT models Pardos and Heffernan (2010); Yudelson et al. (2013), to deep learning KT models based on LSTMs Piech et al. (2015), self-attention Pandey and Karypis (2019); Ghosh et al. (2020) and graphs Yang et al. (2020), to more recent LLM based KT models Wang et al. (2025). Despite advances in the underlying KT models, the performance difference in terms of accuracy is often relatively small Ozyurt et al. (2024) and KT models have low specificity, which results in incorrect answers going undetected Yamkovenko et al. (2025).
In this work we address the issue of incorrect student responses going undetected by considering predictive uncertainty in KT models. We utilise the Monte Carlo Dropout approach Gal and Ghahramani (2016) to estimate predictive uncertainty in four different KT models to compare the impact of predictive uncertainty. We show that the predictive uncertainty of all four models provides a meaningful signal to detect misclassified incorrect student responses, which is pedagogically important for any online learning platform to ensure that student misconceptions do not go undetected.
@@ 2.1 Knowledge Tracing Models
We adapt the standard formulation of knowledge tracing Piech et al. (2015), which focuses on predicting a student’s performance on the next exercise based on their time-series interaction history. Formally, for each student, the sequence of exercises is modeled as $\{e_{i}\}_{i=1}^{t}$, where each exercise $e_{i}$ is represented as a three-tuple: $e_{i}=\{q_{i},c_{i},r_{i}\}$. Here, $q_{i}$ denotes the question ID, $c_{i}$ represents the features associated with $q_{i}$, and $r_{i}\in\{A,B,C,D\}$ is the student’s categorical response corresponding to one of the multi-choice answer options. The features linked to the question may include knowledge component identifiers, question text, explanation text for the answer options, and other pedagogical attributes. The KT model $F_{\theta}$ aims to predict the categorical outcome $\hat{r}_{t+1}$ for the next exercise, conditioned on the history of prior interactions and the information of the upcoming question. The prediction is given by:
@@ 2.1.1 Baseline KT Models
To benchmark KT approaches for mathematics education, we consider four models as instantiations of $F_{\theta}$: Deep Knowledge Tracing (DKT) Piech et al. (2015), Self-Attentive Knowledge Tracing (SAKT) Pandey and Karypis (2019), Attentive Knowledge Tracing (AKT) Ghosh et al. (2020), and our proposed LLM Transformer KT model. DKT models student learning with recurrent neural networks (RNNs). It takes as input the sequence of question identifiers and responses $(q_{i},r_{i})$ and, conditioned on the next question $q_{t+1}$, predicts the probability of a correct response $\hat{r}_{t+1}$ by capturing temporal dependencies in the interaction history. SAKT replaces recurrence with a self-attention mechanism, which allows the model to selectively attend to relevant past interactions when predicting performance on the next question. AKT extends this approach by introducing contextual attention that considers both the similarity between questions and an exponential decay to model student forgetting. Finally, our LLM Transformer KT leverages self-attention over concatenated embeddings, which are generated using the Qwen3 0.6B embedding model Zhang et al. (2025), of the question text and associated text-based question features, and applies cross-attention with the response embeddings.
@@ 2.2 MCDropout for KT Uncertainty
We estimate predictive uncertainty using Monte Carlo Dropout (MC Dropout) Gal and Ghahramani (2016). At inference, dropout remains active and predictions are averaged over $M$ stochastic forward passes. Following the information-theoretic approach based on Shannon entropy (Cover and Thomas, 2006), we quantify total uncertainty by the entropy of the aggregated predictive distribution:
@@
where $\bar{p}_{k}$ denotes the mean predictive probability of class $k$ estimated via MCDropout. Here, $\mathcal{H}_{t}=\{e_{i}\}_{i=1}^{t}$ represents the interaction history of a student up to time $t$, with each exercise $e_{i}=(q_{i},c_{i},r_{i})$ defined by the question identifier $q_{i}$, its associated features or knowledge components $c_{i}$, and the observed response $r_{i}$. Higher entropy corresponds to greater uncertainty in the model’s prediction, whereas lower entropy reflects more confident predictions. In addition to predictive entropy, we compute the standard deviation of the model’s predictions across Monte Carlo samples as a second measure of uncertainty, where further details are provided in Section A.3.
@@ 3 Experiments and Results
We have implemented the proposed approach for four different models. Three of these, DKT Piech et al. (2015), SAKT Pandey and Karypis (2019), and AKT Ghosh et al. (2020), are common baselines often seen in the literature, and the fourth is our proposed self-attention variant that uses an LLM embedding model to process question text rather than a learned embedding of question ids that is more commonly used.
The dataset used to train each of the models contains student responses to questions. Here, we have limited the student response history to contain 100 questions-answer pairs per student. Students answer questions in quiz sessions, where a student will answer 5 questions in one session, which generally increase in difficulty. Therefore, the 100 question-answer pairs will typically be a result of a student working on 20 quizzes. We provide further details on the datasets in Section A.1.
@@ 3.1 Uncertainty Prediction in Knowledge Tracing Models
We also show in Figure 3 that the model predictions have a larger mean entropy when the student chooses an incorrect response. This is largely a result of the models more frequently making incorrect predictions when students choose an incorrect response Yamkovenko et al. (2025). We speculate that the models likely have higher entropy for predictions when students choose an incorrect response due to incorrect responses being a minority class in the dataset and also identifying the specific distractor a student chose is more difficult than deducing if they choose a correct response. This result is further confirmed in Figure 3 where we compare the standard deviation of predictions from each model. In Figure 3 the mean standard deviation of predictions from each model is larger when the student has chosen an incorrect response.
@@ 3.2 How Do Different Models Capture Uncertainty?
We also analyse how uncertainty varies as the student has answered more questions giving the model access to more previous responses. Figure 5 shows how the total entropy varies with question number. Here we can see that DKT has a much higher entropy for the first question than all the other methods. This implies that attention based models become more confident in their initial prediction, which may suggest they can more easily learn a question bias from the training data than the LSTM based models. We also observe a repeating spiking pattern in the entropy value over the question number. This is a result of students answering questions in chunks of five questions per quiz session, which generally increase in difficulty. As a result the more difficult question generates more incorrect responses from students, which we have shown to produce a larger entropy value from the model. We show that this spiking behaviour strongly correlates with the question difficulty in Section C.
@@ 4 Conclusion
Appendix A Experimental Setup
A.1 Dataset Details
@@
A.2 Model Training
We train four knowledge tracing models: DKT Piech et al. (2015), SAKT Pandey and Karypis (2019), AKT Ghosh et al. (2020), and our LLM-based Transformer, under a unified setup. Model sizes are summarized in Table˜2. AKT is the largest (3.3  M parameters), while our LLM Transformer remains compact (2.1  M). Unless otherwise noted, all models share the global training configuration in Table˜3: Adam optimizer, learning rate $3\times 10^{-4}$, cross-entropy loss, batch size 64, and 100 epochs, with a linear warmup followed by a cosine scheduler. Model-specific hyperparameters are listed in Table˜4. We fix embedding and hidden dimensions to 128 and use one layer across models, with dropout tuned per architecture (e.g., 0.5 for SAKT, 0.2 otherwise). The LLM-based model consumes text embeddings truncated to 1024 dimensions, while AKT uses the kq_same setting. All other details are held constant to enable a fair comparison across architectures.
@@ A.2 Model Training
DKT Piech et al. (2015) 1.2 4.890
SAKT Pandey and Karypis (2019) 1.7 6.990
AKT Ghosh et al. (2020) 3.3 13.020
@@ A.3 Model Evaluation
Appendix B Model Quantitative Predictive Accuracy
On the validation split for the multi-class KT task, all models achieve similar accuracy Table˜5. Validation accuracy across training epochs is shown in Figure˜6. Our LLM Transformer attains the highest accuracy ($67.17\%$), narrowly ahead of DKT ($67.07\%$), AKT ($67.13\%$), and SAKT ($67.05\%$). F1 varies more across models: AKT yields the strongest F1 ($35.97\%$), followed by SAKT ($34.15\%$), our model ($33.90\%$), and DKT ($32.57\%$). We also report AUC as a threshold-independent measure of model discrimination. The AUC scores are tightly clustered ranging from 75.35% (DKT) to 76.56% (LLM Transformer). These results suggest that while overall correctness is comparable across architectures, we need to also consider other metrics for AI mathematics tutoring applications.
@@ Appendix B Model Quantitative Predictive Accuracy
DKT Piech et al. (2015) 67.07 32.57 75.35
SAKT Pandey and Karypis (2019) 67.05 34.15 76.22
AKT Ghosh et al. (2020) 67.13 35.97 76.48
Appendix C Uncertainty Analysis
@@
Appendix D Related Works
Knowledge Tracing (KT) models the temporal dynamics of student learning by predicting knowledge states based on sequences of interactions with educational content Corbett and Anderson (2005). Various machine learning methods, including deep neural networks, have been proposed to address this sequential prediction task.
Among these, deep sequential models adopt auto-regressive architectures to capture the progression of student interactions. Deep Knowledge Tracing (DKT) Piech et al. (2015) was one of the first to use an LSTM to estimate students’ knowledge states over time. Subsequent work, such as KQN Lee and Yeung (2019), enhanced KT by encoding student learning activities into knowledge state and skill vectors and modeling their interactions with the dot product.
Memory-augmented models incorporate external memory mechanisms to better capture latent relationships between knowledge concepts (KC) and student knowledge states. DKVMN Zhang et al. (2017) introduced a static key memory for KC representations and a dynamic value memory to track students’ evolving mastery levels.
Graph-based models use graph neural networks to model intrinsic relations among questions, KCs and question-response interactions. Liu et al. (2020) enhanced the KT model by introducing a question-KC bipartite graph to capture intrinsic relationships at both the question and knowledge component levels, thereby augmenting their respective representations. Yang et al. (2021) employed graph convolutional networks to capture interdependencies among KCs and questions more explicitly.
Attention-based models leverage attention mechanisms to dynamically weigh past interactions. SAKT Pandey and Karypis (2019) introduced a self-attention architecture to model the relevance of historical responses. AKT Ghosh et al. (2020) introduced monotonic attention modules to model forgetting behavior.
Despite these advances, uncertainty modeling in KT remains underexplored. Most existing models produce point predictions without quantifying the confidence in knowledge estimates. This is an important limitation for real-world educational applications. UKT Cheng et al. (2025) is the only other work to the best of our knowledge that introduces a KT model to capture uncertainty in student learning by representing knowledge states as stochastic distributions and modeling their transitions with a Wasserstein self-attention mechanism.

### v2

@@ Abstract
The main focus of research on Knowledge Tracing (KT) models is on model developments with the aim of improving predictive accuracy. Most of these models make the most incorrect predictions when students choose a distractor, leading to student errors going undetected. We present an approach to add new capabilities to KT models by capturing predictive uncertainty and demonstrate that a larger predictive uncertainty aligns with model incorrect predictions. We show that uncertainty in KT models is informative and that this signal is pedagogically useful. This is important for an educational learning platform where understanding student ability is necessary.
@@ 1 Introduction
Knowledge Tracing (KT) models are widely used to predict a student’s response to a future question given their interaction history. KT models have developed over time from earlier work on Bayesian KT models [10, 16], to deep learning KT models based on LSTMs [11], self-attention [9, 5] and graphs [14], to more recent LLM based KT models [12]. Despite advances in the underlying KT models, the performance difference in terms of accuracy is often relatively small [8] and KT models have low specificity, which results in incorrect answers going undetected [13].
In this work we address the issue of incorrect student responses going undetected by considering predictive uncertainty in KT models. We utilise the Monte Carlo Dropout approach [4] to estimate predictive uncertainty in four different KT models to compare the impact of predictive uncertainty. We show that the predictive uncertainty of all four models provides a meaningful signal to detect misclassified incorrect student responses, which is pedagogically important for any online learning platform to ensure that student misconceptions do not go undetected.
2 Related Works
Knowledge Tracing (KT) models the temporal dynamics of student learning by predicting knowledge states based on sequences of interactions with educational content [2]. Various machine learning methods, including deep neural networks, have been proposed to address this sequential prediction task.
Among these, deep sequential models adopt auto-regressive architectures to capture the progression of student interactions. Deep Knowledge Tracing (DKT) [11] was one of the first to use an LSTM to estimate students’ knowledge states over time. Subsequent work, such as KQN [6], enhanced KT by encoding student learning activities into knowledge state and skill vectors and modeling their interactions with the dot product.
Memory-augmented models incorporate external memory mechanisms to better capture latent relationships between knowledge concepts (KC) and student knowledge states. DKVMN [17] introduced a static key memory for KC representations and a dynamic value memory to track students’ evolving mastery levels.
Graph-based models use graph neural networks to model intrinsic relations among questions, KCs and question-response interactions. [7] enhanced the KT model by introducing a question-KC bipartite graph to capture intrinsic relationships at both the question and knowledge component levels, thereby augmenting their respective representations. [15] employed graph convolutional networks to capture interdependencies among KCs and questions more explicitly.
Attention-based models leverage attention mechanisms to dynamically weigh past interactions. SAKT [9] introduced a self-attention architecture to model the relevance of historical responses. AKT [5] introduced monotonic attention modules to model forgetting behavior.
Despite these advances, uncertainty modeling in KT remains underexplored. Most existing models produce point predictions without quantifying the confidence in knowledge estimates. This is an important limitation for real-world educational applications. UKT [1] is the only other work to the best of our knowledge that introduces a KT model to capture uncertainty in student learning by representing knowledge states as stochastic distributions and modeling their transitions with a Wasserstein self-attention mechanism.
@@ 3.1 Knowledge Tracing Models
We adapt the standard formulation of knowledge tracing [11], which focuses on predicting a student’s performance on the next exercise based on their time-series interaction history. Formally, for each student, the sequence of exercises is modeled as $\{e_{i}\}_{i=1}^{t}$, where each exercise $e_{i}$ is represented as a three-tuple: $e_{i}=\{q_{i},c_{i},r_{i}\}$. Here, $q_{i}$ denotes the question ID, $c_{i}$ represents the features associated with $q_{i}$, and $r_{i}\in\{A,B,C,D\}$ is the student’s categorical response corresponding to one of the multi-choice answer options. The features linked to the question may include knowledge component identifiers, question text, explanation text for the answer options, and other pedagogical attributes. The KT model $F_{\theta}$ aims to predict the categorical outcome $\hat{r}_{t+1}$ for the next exercise, conditioned on the history of prior interactions and the information of the upcoming question. The prediction is given by:
@@ 3.1.1 Baseline KT Models
To benchmark KT approaches for mathematics education, we consider four models as instantiations of $F_{\theta}$: Deep Knowledge Tracing (DKT) [11], Self-Attentive Knowledge Tracing (SAKT) [9], Attentive Knowledge Tracing (AKT) [5], and our proposed LLM Transformer KT model. DKT models student learning with recurrent neural networks (RNNs). It takes as input the sequence of question identifiers and responses $(q_{i},r_{i})$ and, conditioned on the next question $q_{t+1}$, predicts the probability of a correct response $\hat{r}_{t+1}$ by capturing temporal dependencies in the interaction history. SAKT replaces recurrence with a self-attention mechanism, which allows the model to selectively attend to relevant past interactions when predicting performance on the next question. AKT extends this approach by introducing contextual attention that considers both the similarity between questions and an exponential decay to model student forgetting. Finally, our LLM Transformer KT leverages self-attention over concatenated embeddings, which are generated using the Qwen3 0.6B embedding model [18], of the question text and associated text-based question features, and applies cross-attention with the response embeddings.
@@ 3.2 MCDropout for KT Uncertainty
We estimate predictive uncertainty using Monte Carlo Dropout (MC Dropout) [4]. At inference, dropout remains active and predictions are averaged over $M$ stochastic forward passes. Following the information-theoretic approach based on Shannon entropy [3], we quantify total uncertainty by the entropy of the aggregated predictive distribution:
@@
where $\bar{p}_{k}$ denotes the mean predictive probability of class $k$ estimated via MCDropout. Here, $\mathcal{H}_{t}=\{e_{i}\}_{i=1}^{t}$ represents the interaction history of a student up to time $t$, with each exercise $e_{i}=(q_{i},c_{i},r_{i})$ defined by the question identifier $q_{i}$, its associated features or knowledge components $c_{i}$, and the observed response $r_{i}$. Higher entropy corresponds to greater uncertainty in the model’s prediction, whereas lower entropy reflects more confident predictions. In addition to predictive entropy, we compute the standard deviation of the model’s predictions across Monte Carlo samples as a second measure of uncertainty, where further details are provided in Section 0.A.3.
@@ 4 Experiments and Results
We have implemented the proposed approach for four different models. Three of these, DKT [11], SAKT [9], and AKT [5], are common baselines often seen in the literature, and the fourth is our proposed self-attention variant that uses an LLM embedding model to process question text rather than a learned embedding of question ids that is more commonly used.
The dataset used to train each of the models contains student responses to questions. Here, we have limited the student response history to contain 100 questions-answer pairs per student. Students answer questions in quiz sessions, where a student will answer 5 questions in one session, which generally increase in difficulty. Therefore, the 100 question-answer pairs will typically be a result of a student working on 20 quizzes. We provide further details on the datasets in Section 0.A.1.
@@ 4.1 Uncertainty Prediction in Knowledge Tracing Models
We also show in Figure 3 that the model predictions have a larger mean entropy when the student chooses an incorrect response. This is largely a result of the models more frequently making incorrect predictions when students choose an incorrect response [13]. We speculate that the models likely have higher entropy for predictions when students choose an incorrect response due to incorrect responses being a minority class in the dataset and also identifying the specific distractor a student chose is more difficult than deducing if they choose a correct response. This result is further confirmed in Figure 3 where we compare the standard deviation of predictions from each model. In Figure 3 the mean standard deviation of predictions from each model is larger when the student has chosen an incorrect response.
@@ 4.2 How Do Different Models Capture Uncertainty?
We also analyse how uncertainty varies as the student has answered more questions giving the model access to more previous responses. Figure 5 shows how the total entropy varies with question number. Here we can see that DKT has a much higher entropy for the first question than all the other methods. This implies that attention based models become more confident in their initial prediction, which may suggest they can more easily learn a question bias from the training data than the LSTM based models. We also observe a repeating spiking pattern in the entropy value over the question number. This is a result of students answering questions in chunks of five questions per quiz session, which generally increase in difficulty. As a result the more difficult question generates more incorrect responses from students, which we have shown to produce a larger entropy value from the model. We show that this spiking behaviour strongly correlates with the question difficulty in Section 0.C.
@@ 5 Conclusion
{credits}
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix
Appendix 0.A Experimental Setup
0.A.1 Dataset Details
@@
0.A.2 Model Training
We train four knowledge tracing models: DKT [11], SAKT [9], AKT [5], and our LLM-based Transformer, under a unified setup. Model sizes are summarized in Table˜2. AKT is the largest (3.3  M parameters), while our LLM Transformer remains compact (2.1  M). Unless otherwise noted, all models share the global training configuration in Table˜3: Adam optimizer, learning rate $3\times 10^{-4}$, cross-entropy loss, batch size 64, and 100 epochs, with a linear warmup followed by a cosine scheduler. Model-specific hyperparameters are listed in Table˜4. We fix embedding and hidden dimensions to 128 and use one layer across models, with dropout tuned per architecture (e.g., 0.5 for SAKT, 0.2 otherwise). The LLM-based model consumes text embeddings truncated to 1024 dimensions, while AKT uses the kq_same setting. All other details are held constant to enable a fair comparison across architectures.
@@ 0.A.2 Model Training
DKT [11] 1.2 4.890
SAKT [9] 1.7 6.990
AKT [5] 3.3 13.020
@@ 0.A.3 Model Evaluation
Appendix 0.B Model Quantitative Predictive Accuracy
On the validation split for the multi-class KT task, all models achieve similar accuracy Table˜5. Validation accuracy across training epochs is shown in Figure˜6.
Our LLM Transformer attains the highest accuracy ($67.17\%$), narrowly ahead of DKT ($67.07\%$), AKT ($67.13\%$), and SAKT ($67.05\%$). F1 varies more across models: AKT yields the strongest F1 ($35.97\%$), followed by SAKT ($34.15\%$), our model ($33.90\%$), and DKT ($32.57\%$). We also report AUC as a threshold-independent measure of model discrimination. The AUC scores are tightly clustered ranging from 75.35% (DKT) to 76.56% (LLM Transformer). These results suggest that while overall correctness is comparable across architectures, we need to also consider other metrics for AI mathematics tutoring applications.
@@ Appendix 0.B Model Quantitative Predictive Accuracy
DKT [11] 67.07 32.57 75.35
SAKT [9] 67.05 34.15 76.22
AKT [5] 67.13 35.97 76.48
Appendix 0.C Uncertainty Analysis